OpenStudy (anonymous):

Q is between P and S. R is between Q and S. S is between Q and T. PT=30, QS=16, and PQ=QR=RS. Find PQ. 8 6 16 12

OpenStudy (anonymous):

8. If you draw them on a line from P to T. Then call PQ x, and ST y. Then you get two eqns x+y=14 and 3x+y=30. Solve simult to get x=8. Or equiv. you know that QS=16 and both parts are equal to PQ so 2x=16.

OpenStudy (anonymous):

We know that the distance between Q and S = 16. But we know that QS = QR +RS. And we know taht PQ = QR=Rs. Substiture in we get QS = 16 = 2PQ. hence PQ=16/2 = 8.

OpenStudy (anonymous):

so its 8??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks soooo much :)

OpenStudy (anonymous):

thank yall so much :)